\begin{matrix}T & P \\ T & Q \\ \overline{T} & P & ∧ & Q \end{matrix}

\overline{T} P
\overline{T} P∧Q
\overline{T} Q∧Q
\overline{T} P∧R
T P∧Q
T Q

×

\begin{matrix}\overline{T} & Q \\ \overline{T} & Q & ∧ & Q \end{matrix}

\overline{T} Q ; T P ; T Q ; \overline{T} P∧Q ; ×

\overline{T} Q ; \overline{T} Q ; T P ; T Q ; \overline{T} P∧Q
×

\begin{matrix}T & P \\ T & Q \\ \overline{T} & P & ∧ & Q \end{matrix}

\overline{T} P
\overline{T} P∧Q
\overline{T} Q∧Q
\overline{T} P∧R
T P∧Q
T Q

×

\begin{matrix}\overline{T} & Q \\ \overline{T} & Q & ∧ & Q \end{matrix}

\overline{T} Q ; T P ; T Q ; \overline{T} P∧Q ; ×

\overline{T} Q ; \overline{T} Q ; T P ; T Q ; \overline{T} P∧Q
×